OpenStudy (anonymous):

There are (4^2)^3 ⋅ 4^0 horses on a stud farm. What is the total number of horses on the farm?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[4^0=1\] so this is the same as \[(4^2)^3=4^6\]

OpenStudy (anonymous):

hiiii misty and thank you